OpenStudy (yoyo2443):

Solve and check the inequality, if possible. 6|11m-11|≥66

OpenStudy (mathgeek289):

Btw, I use brainly more, which is why I look lame with 10, but I am actually very good with math. Alright, so we need to solve For M. first, We Must distribute the 6. 6*11M = 66m. 11*6 = 66. So, we now Have: l 66m - 66 l ≥ 66. Now, we must add 66 to the right side. So, we have: 66m ≥ 66+66 66+66 = 120 + 12 = 132 So: 66m ≥ 132. So, we must now divide. 132/66 = 2. So: X ≥ 2

OpenStudy (yoyo2443):

Oh... But wouldn't the 6 that sits outside of the "| |" be subtracted from the 66? making it 6|11m-11|≥66 |11m-11|≥60 ?

OpenStudy (mathgeek289):

No, Because the 6 sitting on the outside is being multiplied by the numbers in the absolute value parenthesizes.

OpenStudy (yoyo2443):

Oh right, makes sense.

OpenStudy (alekos):

I reckon this would reduce to |m - 1| >= 1

OpenStudy (mathgeek289):

No, Because You Must Use The Order Of Operations.

OpenStudy (alekos):

There are two solutions for m The first one you wrote above and the second is m <= 0

OpenStudy (alekos):

The solutions can both be derived from my final equation

OpenStudy (yoyo2443):

how would I check the equation?

OpenStudy (mathgeek289):

Hmmm. I don't think it can be less than zero because the number will stay negative... And negative is less than positive.

OpenStudy (alekos):

If m= -1 Then you would have 6|11(-1)-11| = 6|-11-11| = 6|-22| = 6 × 22 Which is greater than 66

OpenStudy (alekos):

So m can be less than or equal to 0

OpenStudy (alekos):

Do you follow

OpenStudy (yoyo2443):

Yep!